TIFFSV(1) TIFFSV(1) NAME tiffsv−saveanimagefromtheframebufferinaTIFFfile(SiliconGraphicsversion) SYNOPSIS tiffsv[options]output.tif[x1x2y1y2] DESCRIPTION tiffsvsavesallorpartoftheframebufferinafileusingtheTagImageFileFormat,Revision6.0. Bydefault,theimageissavedwithdatasamplespacked(PlanarConfiguration=1),compressed withtheLempel-Ziv&Welchalgorithm(Compression=5),andwitheachstripnomorethan8 kilobytes.Thesecharacteristicscanbeoverriden,orexplicitlyspecifiedwiththeoptionsdescribed below. OPTIONS −bSavetheimageasagreyscaleimageasifitwereprocessedbytiff2bw(1).Thisoptionis includedforcompatibilitywiththestandardscrsave(6D)program. −cSpecifythecompressiontousefordatawrittentotheoutputfile:nonefornocompression,packbitsforPackBitscompression,jpegforbaselineJPEGcompression,zipfor Deflatecompression,andlzwforLempel-Ziv&Welchcompression(default). LZWcompressioncanbespecifiedtogetherwithapredictorvalue.Apredictorvalueof2 causeseachscanlineoftheoutputimagetoundergohorizontaldifferencingbeforeitis encoded;avalueof1forceseachscanlinetobeencodedwithoutdifferencing.LZW-spe-cificoptionsarespecifiedbyappendinga‘‘:’’-separatedlisttothe‘‘lzw’’option;e.g.−c lzw:2forLZWcompressionwithhorizontaldifferencing. −pSpecifytheplanarconfigurationtouseinwritingimagedata.Bydefault,tiffsvwillcreateanewfilewiththedatasamplespackedcontiguously.Specifying−pcontigwill forcedatatobewrittenwithmulti-sampledatapackedtogether,while−pseparatewill forcesamplestobewritteninseparateplanes. −rSpecifythenumberofrows(scanlines)ineachstripofdatawrittentotheoutputfile.By default,tiffsvattemptstosettherows/stripthatnomorethan8kilobytesofdataappear inastrip. NOTE ExceptfortheuseofTIFF,thisprogramisequivalenttothestandardscrsaveprogram.This means,forexample,thatyoucanuseitinconjunctionwiththestandardicutprogramsimplyby creatingalinkcalledscrsave,orbycreatingashellscriptcalledscrsavethatinvokestiffgtwith theappropriateoptions. BUGS Ifdataaresavedcompressedandinseparateplanes,thentherowsineachstripissilentlysetto onetoavoidlimitationsinthelibtiff(3)library. SEEALSO scrsave(6D)pal2rgb(1),tiffdump(1),tiffgt(1),tiffinfo(1),tiffcp(1),tiffmedian(1),libtiff(3) October15,1995 1